# Multiwavelength Observations of the Apparently Non-repeating FRB 20250316A

### v1

@@ Abstract
The physical origin of fast radio bursts (FRBs) remains uncertain. Although multiwavelength observations offer critical diagnostics and have been widely conducted, only Galactic FRB 20200428D is associated with an X-ray burst from the magnetar SGR J1935+2154. Here, we present multiwavelength follow-up observations of the nearby bright FRB 20250316A, including the Five-hundred-meter Aperture Spherical radio Telescope (FAST), Einstein Probe (EP) X-ray mission, Chandra X-ray Observatory, Wide Field Survey Telescope (WFST) and Space Variable Object Monitor/Visible Telescope (SVOM/VT). A 13.08-hour FAST follow-up observational campaign suggests that this burst is likely a one-off event. A prompt EP follow-up and multi-epoch observational campaign totaling 100 ks led to the detection of an X-ray source within the angular resolution of its Follow-up X-ray Telescope (FXT, ). A subsequent Chandra observation revealed this source to be offset by from the FRB position, and established a 0.5-10 keV flux upper limit of at the FRB position, corresponding to at the 40 Mpc distance of the host galaxy NGC 4141. These results set one of the most stringent limits on X-ray emission from a non-repeating FRB, disfavoring ultra-luminous X-ray sources (ULXs) as counterparts of apparently one-off FRBs and offering critical insights into afterglow models. Our study suggests that an arcsecond localization of both the FRB and its potential X-ray counterpart is essential for exploring the X-ray counterpart of an FRB.
@@ 1 Introduction
Fast Radio Bursts (FRBs) are millisecond-duration extragalactic radio transients (D. R. Lorimer et al., 2007). Among the nearly one thousand known FRB sources, fewer than 10% exhibit repeating bursts (L. G. Spitler et al., 2016; CHIME/FRB Collaboration et al., 2021). Although their origin remains unclear, the association between FRB 20200428D and a pair of X-ray pulses from the Milky Way magnetar SGR 1935+2154 (S. Mereghetti et al., 2020; CHIME/FRB Collaboration et al., 2020; C. K. Li et al., 2021) reveals that at least low-luminosity FRBs can be produced by magnetars.
@@
While multiwavelength counterparts of typical extragalactic FRBs have been searched for, no conclusive association has been established (see B. Zhang 2024 for a review). The campaigns include well-studied repeating sources, such as FRB 20121102A (P. Scholz et al., 2016), FRB 20180301A (S. Laha et al., 2022), FRB 20180916B (M. Pilia et al., 2020; P. Scholz et al., 2020), FRB 20190520B (J. Sydnor et al., 2025), FRB 20220912A (A. M. Cook et al., 2024), and FRB 20240114A (F. Eppel et al., 2025). The most stringent X-ray constraints come from the nearby (3.6 Mpc) repeating FRB 20200120E (A. B. Pearlman et al., 2025), which exclude X-ray counterparts similar to ultraluminous X-ray sources (ULXs) or the Crab Nebula. However, most bursts from FRB 20200120E are fainter than typical FRBs (M. Bhardwaj et al., 2021; F. Kirsten et al., 2022; B. Zhang, 2024), suggesting that its multiwavelength signals may differ from those of typical FRBs. For apparently non-repeating FRBs, constraints are usually much looser, because they rely primarily on telescopes with large fields of view (FOVs) and relatively poor sensitivities, such as AGILE (F. Verrecchia et al., 2021), Konus-Wind (A. Ridnaia et al., 2024; G. Waratkar et al., 2025), and Fermi (G. Principe et al., 2023). Although two apparently non-repeating FRBs, FRB 20190608B and FRB 20200430A, have relatively deep constraints, the observations occurred nearly two years after the bursts, and both sources are distant, with redshifts (T. Eftekhari et al., 2023).
@@
FRB 20250316A was detected by the Canadian Hydrogen Intensity Mapping Experiment (CHIME, M. Ng & CHIME/FRB Collaboration 2025; FRB Collaboration et al. 2025a) at 08:33:50.859038(3) on March 16, 2025 (topocentric at 400 MHz), with a dispersion measure (DM) of pc cm-3. The burst is very bright, with a peak flux of kJy and a fluence of kJy ms. The initial localization was (RA, Dec) = (182.476∘, +58.8494∘), with arcminute uncertainties. Subsequent localization using the CHIME core and the three outriggers – KKO near Penticton, the Green Bank Observatory (GBO), and the Hat Creek Radio Observatory (HCO) – refined its position to (RA, Dec) = (12h09m44.319s, +58d50m56.708s), with 1 positional uncertainties of semi-major axis = 68 mas, semi-minor axis = 57 mas at a position angle of east of north (S. Andrew & Chime/Frb Collaboration, 2025; FRB Collaboration et al., 2025a). This FRB is positionally associated with the nearby spiral galaxy NGC 4141, at a distance of Mpc based on the Tully-Fisher relation. An infrared counterpart candidate from James Webb Space Telescope observations was reported (P. K. Blanchard et al., 2025). Assuming a host galaxy distance of 40 Mpc and a central frequency of 600 MHz for CHIME, the isotropic energy and luminosity of the FRB are and , respectively.
@@ 2 Observations
This section provides an overview of our multiwavelength observations of FRB 20250316A in the radio, X-ray, and optical bands. The timeline of these observations relative to the FRB is shown in Figure 1.
@@ 2.1 Radio observations
Following the report of FRB 20250316A (M. Ng & CHIME/FRB Collaboration, 2025), we triggered the Five-hundred-meter Aperture Spherical radio Telescope (FAST) for follow-up monitoring. Eight sessions were carried out between 2025 March 17 and 24 as part of FAST Director Discretion Time and the FAST FRB key science project (ATEL 17126). The observations were conducted in “tracking” mode. During the observations, data from all 19 beams of the FAST L-band receiver ( GHz) were recorded. The data were channelized into 4096 frequency channels and sampled with a time resolution of 49.153 s. Dual linear polarization signals were 8-bit sampled, processed using the Reconfigurable Open Architecture Computing Hardware generation 2 (ROACH 2) and stored in PSRFITS search mode format. A 1 K equivalent noise-switched calibration signal was recorded before each observation for calibration. The total exposure time amounted to 13.08 hours (13 hours and 5 minutes) across all sessions, as detailed in Table 1.
@@
a SNR 7. A burst width of 1 ms was assumed.
@@ 2.2.1 Einstein Probe
The Einstein Probe (EP, W. Yuan et al. 2022) is an X-ray mission launched in January 2024. It carries two instruments: the Wide-field X-ray Telescope (WXT), a monitoring instrument with a field of view of 3600 deg2, and the Follow-up X-ray Telescope (FXT), which provides a large effective area and an angular resolution of . WXT covered the position of FRB 20250316A six times within a 24-hour window centered on the burst, with a total exposure of 2 hours (Figure 1 and Table 2). However, the FRB position was not covered at the burst time. The temporally closest observation started on 2025-03-16 at 07:11:07 UTC and ended at 07:30:49 UTC, about 1 hour before the burst. This was followed by the first post-burst observation, starting on 2025-03-16 at 11:57:59 UTC, about 3.5 hours after the burst, with an exposure time of 1324 s.
Following the discovery of FRB 20250316A, we observed the FRB field using the FXT (Y. Chen et al., 2021) onboard EP. Our campaign consisted of ten observations: nine were taken between 32 hours and 11 days post-burst (Figure 1 and Table 2). An additional late-time observation was carried out on 2025 May 11 to investigate the variability of a previously reported X-ray source in the field, EP J120944.2+585060 (H. Sun et al., 2025a, b).
@@ 2.3.2 SVOM/VT
The Space-based multi-band astronomical Variable Objects Monitor (SVOM) is a China-France joint mission designed to detect gamma-ray bursts (J. Wei et al., 2016). In addition to gamma ray instruments – Gamma-ray Burst Monitor (GRM, gamma-rays), ECLAIR (hard X-rays), and Microchannel X-ray Telescope (MXT, soft X-rays) – SVOM also carries a Visible Telescope (VT). VT is a 43-cm aperture telescope equipped with a dichroic beam splitter, which divides the light into two channels: a blue channel (VTB) with a wavelength range from 0.4 to 0.65 m and a red channel (VTR) from 0.65 to 1 m. The 3 limiting magnitude for a 300-second exposure time is about 22.5 mag. SVOM/VT conducted two runs of observations targeting the field of FRB 20250316A, simultaneously acquiring data in both the VTB and VTR bands. The first run began at 18:22:00 UTC on March 23, 2025. The second run started at 04:22:05 UTC on March 25, 2025. Each exposure lasted 60 seconds. After filtering the images to remove those affected by background noise or distorted stellar PSFs, 101 and 97 good-quality images were obtained in the two runs, respectively, for each band. Detailed information about the observations is provided in Table 3 and Table LABEL:tb:opt1img.
@@ 3.1 Radio
Single pulses were searched in the FAST data using PRESTO (S. Ransom, 2011), with a DM range of 140-180 pc cm-3 and a step size of 0.1 pc cm-3. No pulses with S/N7 were detected, which corresponds to a fluence threshold of 15.3 mJy ms for a fiducial burst width of 1 ms and a bandwidth of 400 MHz (Table 1). This fluence limit translates to an energy upper limit of erg, assuming a source distance of 40 Mpc.
The absence of subsequent bursts, despite exceptionally deep follow-up observations (with a fluence limit of 15.3 mJy ms), combined with the source’s proximity, suggests that FRB 20250316A is an apparently one-off event. It thus represents one of the nearest apparently one-off FRBs with sub-arcsecond localization, making it an excellent target for multiwavelength follow-up studies.
@@ 3.2.1 Einstein Probe
Using the WXT data reduction pipeline and calibration database (CALDB, H. Cheng et al., 2025), we analyzed the observations obtained 1.5 hours before and 3.5 hours after FRB 20250316A. No significant X-ray sources (above a 5 significance threshold) were detected within the positional uncertainty radius ( 3 arcmin) centered at the FRB location (Figure 2, left panel). For these two observations, we extracted the source counts from a circular region with a radius of 9 arcmin at the FRB position, and the background counts from an annulus with inner and outer radii of 18 arcmin and 36 arcmin, as shown in Figure 2. We derived the 90% confidence count rate upper limits of 0.0057 and 0.0036 , respectively, in the 0.5–4 keV range. Assuming an absorbed power-law spectrum with a photon index of 2 and a Galactic absorption column density of , the flux limits are and , respectively, corresponding to and assuming a luminosity distance of 40 Mpc.
In the first FXT observation, which began at 2025-03-17T16:57:29 with an exposure time of 2.9 ks, no X-ray sources were detected within the CHIME error region. In the second observation, starting at 2025-03-18T15:23:50 with the same exposure time of 2.9 ks, FXT detected a weak, uncatalogued X-ray source, designated as EP J120944.2+585060. The position of this source is generally consistent with the refined localization provided by CHIME (C. Leung & Chime/Frb Collaboration, 2025). The source was observed in all eight subsequent FXT observations. For each observation, we extracted the source using a circular region with a 1-arcmin radius. The background was selected from an annulus with inner and outer radii of 1 arcmin and 3 arcmin, where we excluded contamination from bright sources. The FXT data were processed using the FXT data analysis software (fxtsoftware v1.20; http://epfxt.ihep.ac.cn/analysis). The process involved particle event identification, pulse invariant conversion, grade calculation and selection (grade 0–12), bad and hot pixel flag and selection of good time intervals using housekeeping files.
The spectrum extracted from the stacked images of all the FXT observations (Figure 2, middle panel) was fitted in XSPEC with an absorbed power-law model. The best-fit photon index is (uncertainties at the confidence level), assuming the same Galactic absorption as previously introduced. When modeling the spectrum with the absorption as a free parameter, we found negligible intrinsic absorption. The photon index remains consistent with that obtained from the fixed-absorption fit.
@@ 3.2.2 Chandra
We used the Chandra Interactive Analysis of Observations (CIAO, A. Fruscione et al. 2006; version 4.17) software package and the Chandra Calibration Database (CalDB; version 4.12) to reduce and analyze the Chandra observations. By applying the vtpdetect in CIAO for source detection, we identified an X-ray source (CXOU J120943.7+585051) with a significance of 8 within the FXT error circle (Figure 2, right panel). The source is located at RA (J2000) = 12:09:43.72, Dec (J2000) = +58:50:51.4, with a positional uncertainty of 0.74 arcsec (68% confidence level, statistical + systematic). Assuming an absorbed power-law model with a photon index and a hydrogen column density , we estimated the unabsorbed 0.5–10 keV flux from the count rate using PIMMS (K. Mukai, 1993). The resulting flux is (1 uncertainties), generally consistent with previous EP-FXT measurements, thereby confirming this source as EP J120944.2+585060. The position of this source is offset by 7 arcseconds from the precise localization of FRB 20250316A provided by the full CHIME/FRB Outriggers (FRB Collaboration et al., 2025b). Considering the sub-arcsecond uncertainties of both the FRB and Chandra/HRC positions, this relatively large spatial separation suggests that EP J120944.2+585060 is not physically associated with FRB 20250316A.
@@
We note the presence of another object, CXOU J120946.7+585054 (green circle in the right panel of Figure 2), at a 3 sigma confidence level within 1 arcmin of the FRB. It is away from CXOU J120943.7+585051 and away from the nominal EP J120944.2+585060. It is difficult for FXT to resolve these two objects, as shown in the stacked FXT image (middle panel of Figure 2). Thus, the fluxes of CXOU J120943.7+585051 estimated from the FXT observations are contaminated by 30% due to CXOU J120946.7+585054. This contamination may explain why the FXT fluxes are slightly higher than the Chandra flux, although the difference is not statistically significant.
@@
At the position of FRB 20250316A, we derive a 90% upper limit on the unabsorbed 0.5 – 10 keV flux of , assuming an absorbed power-law spectrum with a photon index of 2 and . For a distance of 40 Mpc from the host galaxy, this corresponds to an X-ray luminosity upper limit of .

### v2

@@ Abstract
The physical origin of fast radio bursts (FRBs) remains uncertain. Although multiwavelength observations have been widely conducted, only Galactic FRB 20200428D is associated with an X-ray burst from the magnetar SGR J1935+2154. Here, we present multiwavelength follow-up observations of the nearby bright FRB 20250316A, including the Five-hundred-meter Aperture Spherical radio Telescope (FAST), Einstein Probe (EP) X-ray mission, Chandra X-ray Observatory, Wide Field Survey Telescope (WFST) and Space Variable Object Monitor/Visible Telescope (SVOM/VT). The 13.08-hour FAST follow-up campaign without pulse detection requires an energy distribution flatter than those of well-known repeating FRBs, suggesting that this burst is likely a one-off event. A prompt EP follow-up and multi-epoch observational campaign totaling 100 ks led to the detection of an X-ray source within the angular resolution of its Follow-up X-ray Telescope (FXT, ). A subsequent Chandra observation revealed this source to be offset by from the FRB position, and established a 0.5-10 keV flux upper limit of at the FRB position, corresponding to at the 40 Mpc distance of the host galaxy NGC 4141. These results set one of the most stringent limits on X-ray emission from a non-repeating FRB, disfavoring ultra-luminous X-ray sources (ULXs) as counterparts of apparently one-off FRBs and offering critical insights into afterglow models. Our study suggests that an arcsecond localization of both the FRB and its potential X-ray counterpart is essential for exploring the X-ray counterpart of an FRB.
@@ 1 Introduction
Fast Radio Bursts (FRBs) are millisecond-duration extragalactic radio transients (D. R. Lorimer et al., 2007). Among the nearly one thousand known FRB sources, fewer than 10% exhibit repeating bursts (L. G. Spitler et al., 2016; CHIME/FRB Collaboration et al., 2021; Chime/Frb Collaboration et al., 2023a). Although their origin remains unclear, the association between FRB 20200428D and a pair of X-ray pulses from the Milky Way magnetar SGR 1935+2154 (S. Mereghetti et al., 2020; CHIME/FRB Collaboration et al., 2020; C. K. Li et al., 2021) reveals that at least low-luminosity FRBs can be produced by magnetars.
@@
While multiwavelength counterparts of typical extragalactic FRBs have been searched for, no conclusive association has been established (see B. Zhang 2024 for a review). The campaigns include well-studied repeating sources, such as FRB 20121102A (P. Scholz et al., 2016), FRB 20180301A (S. Laha et al., 2022), FRB 20180916B (M. Pilia et al., 2020; P. Scholz et al., 2020), FRB 20190520B (J. Sydnor et al., 2025), FRB 20220912A (A. M. Cook et al., 2024), and FRB 20240114A (F. Eppel et al., 2025). The most stringent limits in X-ray luminosity come from the nearby (3.6 Mpc) repeating FRB 20200120E (A. B. Pearlman et al., 2025), which exclude X-ray counterparts similar to ultraluminous X-ray sources (ULXs) or the Crab Nebula. However, most bursts from FRB 20200120E are fainter than typical FRBs (M. Bhardwaj et al., 2021; F. Kirsten et al., 2022; B. Zhang, 2024), and its multiwavelength signals may also differ from those of typical FRBs. For apparently non-repeating FRBs, constraints are usually much looser, because they rely primarily on telescopes with large fields of view (FOVs) and relatively poor sensitivities, such as AGILE (F. Verrecchia et al., 2021), Konus-Wind (A. Ridnaia et al., 2024), AstroSat-CZTI (G. Waratkar et al., 2025), and Fermi (G. Principe et al., 2023). Although two apparently non-repeating FRBs, FRB 20190608B and FRB 20200430A, constrain X-ray luminosities to be , the observations occurred nearly two years after the bursts, and both sources are distant, with redshifts (T. Eftekhari et al., 2023).
@@
FRB 20250316A was detected by the Canadian Hydrogen Intensity Mapping Experiment (CHIME, M. Ng & CHIME/FRB Collaboration 2025; CHIME/FRB Collaboration et al. 2025) at 08:33:50.859038(3) UTC on March 16, 2025 (topocentric at 400 MHz). The burst is very bright, with a peak flux of kJy and a fluence of kJy ms. High-cadence radio follow-up radio observations detected no pulse (O. S. Ould-Boukattine et al., 2025; CHIME/FRB Collaboration et al., 2025). The initial localization has arcminute uncertainties. Subsequent localization using the CHIME core and its three outriggers refined the position to (RA, Dec) = (12h09m44.319s, +58d50m56.708s), with 1 positional uncertainties of 68 mas (semi-major axis) and 57 mas (semi-minor axis) at a position angle of east of north (S. Andrew & Chime/FRB Collaboration, 2025; CHIME/FRB Collaboration et al., 2025). This FRB is positionally associated with the nearby spiral galaxy NGC 4141, at a distance of Mpc based on the Tully-Fisher relation(J. G. Sorce et al., 2014; R. B. Tully et al., 2016). While an infrared counterpart candidate from James Webb Space Telescope observations was reported (P. K. Blanchard et al., 2025a), no clear optical counterpart was detected in MMT and Gemini observations (CHIME/FRB Collaboration et al., 2025). Assuming a host galaxy distance of 40 Mpc and a central frequency of 600 MHz for CHIME, the isotropic energy and luminosity of the FRB are and , respectively.
@@ 2 Observations
This section summarizes our multiwavelength observations of FRB 20250316A in radio, X-ray, and optical bands. The timeline relative to the FRB is shown in Figure 1.
@@ 2.1 Radio observations
a A burst width of 1 ms was assumed. SNR 7.
Following the report of FRB 20250316A (M. Ng & CHIME/FRB Collaboration, 2025), we triggered the Five-hundred-meter Aperture Spherical radio Telescope (FAST) for follow-up monitoring. Eight sessions between 2025 March 17 and 24 were conducted as part of FAST Director Discretion Time and the FAST FRB key science project (L. Qian et al., 2025) in tracking mode. Data from all 19 beams of the FAST L-band receiver (– GHz) were channelized into 4096 frequency channels, sampled at 49.153 s, and recorded in PSRFITS search mode with dual linear polarization signals 8-bit sampled and processed using ROACH 2. A 1 K noise-switched calibration signal was taken before each observation. The total exposure time was 13.08 hour (13 hours and 5 minutes) across all sessions (Table 1).
@@ 2.2.1 Einstein Probe
The Einstein Probe (EP, W. Yuan et al. 2022) is an X-ray mission launched in January 2024. It carries two instruments: the Wide-field X-ray Telescope (WXT), a monitoring instrument with a field of view of 3600 deg2, and the Follow-up X-ray Telescope (FXT), which provides a large effective area and an angular resolution of . WXT covered the position of FRB 20250316A in the survey mode six times within a 24-hour window centered on the burst, with a total exposure of 2 hours (Figure 1 and Table 2). However, the FRB position was not covered at the burst time. The temporally closest observation started on 2025-03-16 at 07:11:07 UTC and ended at 07:30:49 UTC, about 1 hour before the burst. This was followed by the first post-burst observation, starting on 2025-03-16 at 11:57:59 UTC, about 3.5 hours after the burst, with an exposure time of 1324 s.
Following the discovery of FRB 20250316A, we observed the FRB field using the FXT (Y. Chen et al., 2021) onboard EP. Our campaign consisted of ten target-of-opportunity observations: nine were taken between 32 hours and 11 days post-burst (Figure 1 and Table 2). An additional late-time observation was carried out on 2025 May 11 to investigate the variability of a previously reported X-ray source in the field, EP J120944.2+585060 (H. Sun et al., 2025a, b).
@@ 2.3.2 SVOM/VT
The Space-based multi-band astronomical Variable Objects Monitor (SVOM) is a China–France mission designed to detect gamma-ray bursts (J. Wei et al., 2016). In addition to gamma-ray instruments—GRM (gamma-rays), ECLAIR (hard X-rays), and MXT (soft X-rays)—SVOM carries a 43-cm Visible Telescope (VT) with a dichroic beam splitter dividing light into two channels: VT_B (0.4–0.65 m) and VT_R (0.65–1 m). SVOM/VT conduected two runs of observations targeting FRB 20250316A, simultaneously acquiring VT_B and VT_R data. The first run began 2025-03-23 18:22:00 UTC, the second on 2025-03-25 04:22:05 UTC. Each exposure was 60 s. After removing images affected by background noise or distorted PSFs, 101 and 97 good-quality frames were obtained in the two runs for each band. Details are listed in Tables 3 and LABEL:tb:opt1img.
@@ 3.1 Radio
Single pulses were searched for in the FAST data using PRESTO (S. M. Ransom, 2001; S. M. Ransom et al., 2002, 2003), with a DM range of 140-180 pc cm-3 and a step size of 0.1 pc cm-3. No pulses with S/N7 were detected, which corresponds to a fluence threshold of 15.3 mJy ms for a fiducial burst width of 1 ms and a bandwidth of 400 MHz (Table 1). This fluence limit translates to an energy upper limit of erg, assuming a source distance of 40 Mpc.
@@ 3.2.1 Einstein Probe
Using the WXT data reduction pipeline and calibration database (CALDB, H. Cheng et al., 2025), we analyzed the observations obtained 1.5 hours before and 3.5 hours after FRB 20250316A. No significant X-ray sources (above a 5 significance threshold) were detected within the positional uncertainty radius ( 3 arcmin) centered at the FRB location (Figure 2, upper left panel). For these two observations, we extracted the source counts from a circular region with a radius of 9 arcmin at the FRB position, enclosing over 90% of the photons within the focal spot region as determined by the WXT ground calibration, and the background counts from an annulus with inner and outer radii of 18 arcmin and 36 arcmin, respectively, as shown in Figure 2. We derived the 90% confidence count rate upper limits of 0.0057 counts/s and 0.0036 counts/s, respectively, in the 0.5–4 keV range. Assuming an absorbed power-law spectrum with a photon index of 2 and a Galactic absorption column density of , the flux limits are and , respectively, corresponding to and assuming a luminosity distance of 40 Mpc.
In the first FXT observation, which began at 2025-03-17T16:57:29 with an exposure time of 2.9 ks, no X-ray sources were detected within the CHIME error region. In the second observation, starting at 2025-03-18T15:23:50 with the same exposure time of 2.9 ks, FXT detected a weak, uncatalogued X-ray source, designated as EP J120944.2+585060. The position of this source is generally consistent with the first refined localization provided by CHIME (C. Leung & Chime/Frb Collaboration, 2025). The source was observed in all eight subsequent FXT observations. For each observation, we extracted the source using a circular region with a 1-arcmin radius. The background was selected from an annulus with inner and outer radii of 1 arcmin and 3 arcmin, respectively, where we excluded contamination from bright sources. The FXT data were processed using the FXT data analysis software (fxtsoftware v1.20; http://epfxt.ihep.ac.cn/analysis). The process involved particle event identification, pulse invariant conversion, grade calculation and selection (grade 0–12), bad and hot pixel flag and selection of good time intervals using housekeeping files.
The spectrum extracted from the stacked images of all the FXT observations (Figure 2, upper right panel) was fitted in XSPEC (v12.14.1) with an absorbed power-law model. The best-fit photon index is (uncertainties at the confidence level), assuming the same Galactic absorption as previously introduced. When modeling the spectrum with the absorption as a free parameter, we found negligible intrinsic absorption. The photon index remains consistent with that obtained from the fixed-absorption fit.
@@ 3.2.2 Chandra
We used the Chandra Interactive Analysis of Observations (CIAO, A. Fruscione et al. 2006; version 4.17) software package and the Chandra Calibration Database (CalDB; version 4.12) to reduce and analyze the Chandra observations. By applying the vtpdetect in CIAO for source detection, we identified an X-ray source (CXOU J120943.7+585051) within the FXT error circle (Figure 2, lower left panel). The source is located at RA (J2000) = 12:09:43.72, Dec (J2000) = +58:50:51.4, with a positional uncertainty of 0.74 arcsec (68% confidence level, statistical + systematic). Employing a circular source region with a radius of 1 arcsec and a background annulus with inner and outer radii of 3 arcsec and 5 arcsec, respectively, we obtained a detection significance of 8 using the Li-Ma formula (T. P. Li & Y. Q. Ma, 1983). The source count rate was derived using the srcflux tool, yielding a measurement of counts/s (1 uncertainties, 0.1–10.0 keV). Assuming an absorbed power-law model with a photon index and a hydrogen column density , we estimated the unabsorbed 0.5–10 keV flux from the count rate using PIMMS (K. Mukai, 1993) using the average spectral response files provided within the tool. The resulting flux is (1 uncertainties), generally consistent with previous EP-FXT measurements, thereby confirming this source as EP J120944.2+585060. The position of this source is offset by 7 arcseconds from the precise localization of FRB 20250316A provided by the full CHIME/FRB Outriggers (FRB Collaboration et al., 2025). The combined positional uncertainty for the two sources is approximately 0.7 arcsecond (derived from the sub-arcsecond uncertainties of both the FRB and Chandra/HRC positions), meaning the observed spatial separation corresponds to nearly 10. This relatively large spatial separation suggests that EP J120944.2+585060 is not physically associated with FRB 20250316A.
@@
We note the presence of another object, CXOU J120946.7+585054, detected at 4 confidence level within 1 arcmin of the FRB, at the coordinates of RA (J2000) = 12:09:46.71, Dec (J2000) = +58:50:54.1 (with a positional uncertainty of 0.93 arcsec, 68% confidence level, statistical + systematic, green circle in the lower left panel of Figure 2). It is away from CXOU J120943.7+585051 and away from the nominal EP J120944.2+585060. It is difficult for FXT to resolve these two objects, as shown in the stacked FXT image (upper right panel of Figure 2). The measured count rate for CXOU J120946.7+585054, obtained with srcflux, is counts/s (1 uncertainties, 0.1–10.0 keV). Assuming an absorbed power-law model with a photon index and the galactic absorption, we estimated the unabsorbed 0.5–10 keV flux from the count rate using PIMMS to be . Thus, the flux of CXOU J120943.7+585051 measured from the FXT observations are contaminated by 30% due to CXOU J120946.7+585054. This contamination may explain why the FXT fluxes are slightly higher than the Chandra flux, although the difference is not statistically significant.
@@
At the position of FRB 20250316A, we derive a 90% upper limit on the unabsorbed 0.5 – 10 keV flux of , assuming an absorbed power-law spectrum with a photon index of 2 and . For a distance of 40 Mpc from the host galaxy, this corresponds to a 0.5-10 keV X-ray luminosity upper limit of .
3.2.3 Properties of CXOU J120943.7+585051
Although the precise location of CXOU J120943.7+585051/EP J120944.2+585060 provided by Chandra rules out its association with FRB 20250316A, the nature of the X-ray source remains of interest. The source was detected by FXT in observations spanning over 50 days. The keV light curve extracted at the Chandra position shows moderate variability (see Table 2 and Figure 3). However, a test against a constant flux model yielded for 8 degrees of freedom (dof), corresponding to a p-value of 0.419. This high p-value indicates that the observed data are consistent with the null hypothesis of no variability, showing no statistically significant deviation from a constant flux. Together, these results indicate that the flux remained stable within uncertainties across the earlier observations. The late-time observation in May further confirms this steady behavior.
With an estimated flux of inferred from the Chandra observation, this source, if associated with the host galaxy at 40 Mpc, corresponds to a luminosity of . This places it within the typical luminosity range for ULXs, which are thought to be powered by accretion onto a compact object (either a stellar-mass black hole, neutron star, or potentially an intermediate-mass black hole) at a luminosity comparable to or above the Eddington luminosity of stellar black holes (see P. Kaaret et al. 2017 for a review). Alternatively, if CXOU J120943.7+585051 is not associated with the FRB’s host galaxy, it could be a foreground Galactic source (such as a cataclysmic variable) or a background extragalactic object like an active galactic nucleus (AGN). Distinguishing between these possibilities would require multiwavelength observations.